# Bio-Inspired Classification: Combining Information Theory and Spiking Neural Networks - Influence of the Learning Rules

### v1

@@ II-C4 SpikeProp Algorithm
II-C5 Chronotron Learning Rule
The Chronotron minimizes the actual and target spike times [16, 17]
@@ II-C5 Chronotron Learning Rule
II-C6 Remote Supervised Method
@@
II-D Hybrid Learning Algorithms
Hybrid algorithms integrate elements of both supervised and unsupervised learning.
II-D1 Artificial-to-Spiking Neural Network conversion
@@
II-D2 Reward-based STDP
Reward-based STDP modifies STDP with reinforcement learning [20]
@@ II-D2 Reward-based STDP
II-D3 Bio-inspired Active Learning
Active learning updates weights based on uncertainty [37, 38, 36]
@@ III Classification Task
The SNN neural network under consideration consists of three layers: input, hidden, and output, each containing $n$ neurons; see Figure 1. It was based on the Leakly-Integrate-and-Fire (LIF) biophysical neuron model that is a fundamental component in SNNs for simulating the subthreshold behavior of a neuron’s membrane potential $U(t)\in\mathbb{R}$ [40, 39]. Given an input vector $\mathbf{x}(t)\in\mathbb{R}^{n}$, synaptic weights $\mathbf{w}\in\mathbb{R}^{n}$, and bias $b\in\mathbb{R}$, the net input current is defined as:
@@
We considered different sets of neuron parameters and the presented results include the most efficient set of parameters for each learning algorithm. The threshold took values 1.00, 0.50, 0.40, 0.30, 0.20, 0.10, and 0.05. The decline was 0.100, 0.05, 0.03, 0.01, and the learning rates 0.0500, 0.0100, 0.0010, 0.0098, and 0.0001. We considered neural networks composed of 16, 32, 64, 128, 256, 512, and 1024 neurons in layers. The best of the obtained results are presented in Table II, Table III, and Table IV.
@@
where $\alpha=2$ for binary sequences. For purely random sequences, $c_{\alpha}(\textbf{x}_{n}^{1})$ converges to 1, while for deterministic sequences, it approaches 0. This measure effectively approximates the entropy of ergodic stochastic processes, making it a valuable tool to evaluate randomness and structure within the output of the neural network [41, 42, 43]. For completeness, we performed our analysis for information sources for which the subsequent symbols are more and more correlated [44]. Thus, we consider three types of dataset that consist of Bernoulli sequences (uncorrelated), two-state Markov, and Poisson processes. The Poisson process has been often used to model spike trains since early studies, due to the proportionality between mean and variance in many neural responses. Experimental data have confirmed this Poisson property [45]. The binary sequence of each process is converted into inputs to LIF models with specified $\Delta t=1ms$ [11]. For the evaluation of the proposed classification task, we applied accuracy, the mean squared error (MSE), and the mean absolute error (MEA).
IV Results and Discussion
@@
We consider classical and bio-inspired learning algorithms in SNNs. Table II presents the influence of the unsupervised learning algorithm such as Hebbian, STDP, and SDSP on the performance of SNNs. All computation were provided during 10 epochs, on Intel(R) Core(TM) i7-14700F, 2.10 GHz. A general finding of our study is the pronounced dependence of both classification accuracy and computational overhead on the choice of learning algorithm and the underlying characteristics of the input data. Observed classification accuracies ranged from 89.50% to 100.00% across the evaluated methods, whereas computational times exhibited substantial variation, ranging from approximately 7 seconds to over 48,000 seconds.
@@
Table II presents the influence of the unsupervised learning algorithm such as Hebbian, STDP, and SDSP on the performance of SNNs. The results obtained for the Hebbian algorithm demonstrate that, within the classical framework, high classification accuracy was achieved for the Bernoulli and Markov processes, with values of 99.00% and 95.79%, respectively. These results were accompanied by relatively long computation times of 52.93 s and 74.66 s. Furthermore, MSE values of 0.0011 and 0.00422, along with MAE values of 0.0100 and 0.0422, respectively, suggest minimal prediction errors. The coefficient of determination $R^{2}$ scores of 0.9600 and 0.8313 further confirm a strong agreement between predicted and actual values. For the Poisson process, a classification accuracy of 90.00% was observed, with a $R^{2}$ score of 0.9074. The corresponding MSE and MAE values were 0.1000, while the $R^{2}$ score of 0.5998 indicates a moderate correlation between the predictions and the actual values. It is worth emphasizing that the accuracy for the Poisson process is achieved for a twice smaller number of neurons in the layers (for Bernoulli and Markov processes, we achieve similar accuracy with the same number of neurons in the layers). An increase in the number of neurons causes a decrease in accuracy below 50.00%, contrary to the case of Bernoulli and Markov processes, where accuracy increases with increasing number of neurons in layers. Thus, the worse classification results for the Poisson process may be related to the fact that events occur randomly in time, making it difficult to generalize Hebbian learning effectively. The relatively low computation time indicates that the Hebbian learning algorithm efficiently processes structured data. However, taking into account the nature of Hebbian learning it is effective when there are strong, predictable relationships between inputs and outputs, like in the case in Bernoulli or Markov processes.
@@
The introduction of other bioinspired learning algorithms, namely STDP and SDSP provides a similar level of accuracy, in a similar computational time, with the exception of the SDSP algorithm for Markov processes, where the computational time is an order of magnitude longer. However, the lower MSE, and MAE for Bernoulli and Markov processes indicate that STDP fine-tunes synaptic weights more effectively than Hebbian learning, reducing classification errors.
@@
TABLE II: Influence of unsupervised learning algorithm on SNN.
Type Subtype Datasets Bio-inspired Epochs $n$ Time [s] Accuracy [%] MSE MAE $R^{2}$ score
Unsupervised Hebbian Bernoulli Yes 10 128 51.93 99.00 0.0100 0.0100 0.9600
Markov Yes 10 128 74.66 95.78 0.0422 0.0422 0.8313
Poisson Yes 10 64 17.17 90.00 0.1000 0.1000 0.5998
STDP Bernoulli Yes 10 64 62.08 99.50 0.0050 0.0050 0.9800
Markov Yes 10 128 513.73 98.00 0.0200 0.0200 0.9200
Poisson Yes 10 64 63.36 89.50 0.1050 0.1050 0.5798
SDSP Bernoulli Yes 10 128 43.2 98.00 0.0200 0.2000 0.9200
Markov Yes 10 128 46.1 93.87 0.0613 0.0613 0.7549
Poisson Yes 10 64 17.83 92.50 0.0750 0.0750 0.6999
Table III presents the influence of the superseded learning algorithm such as backpropagation, tempotron, SpikeProp, Chronotron and ReSuMe on the performance of SNNs. The BP algorithm achieves perfect classification accuracy for Bernoulli, Markov, and Poisson processes, as evidenced by zero MSE and MAE, indicating no classification errors. The $R^{2}$ score of 1.000 further confirms its ability to perfectly predict sequence outputs. However, this unmatched accuracy comes at a significant computational cost, with processing times, making BP the most resource-intensive supervised method.
Although BP is theoretically optimal in terms of predictive performance, its high computational demands render it impractical for real-time or large-scale applications where efficiency is crucial. This trend also holds for the STBP algorithm, but with a significantly greater time expenditure than BP. In turn, in [11], the authors state that they studied Bernoulli sources; however, in their network, the input layer consisted of LIF neurons, while the remaining layers comprised perceptrons trained using the tempotron algorithm. They did not present results for this datasets. We achive the high accuracy of clasiffication, reaching 99.00% for Bernoulli process and 100.00% for Markov process, while in the case of Poisson process hte level of accuracy in lower, namely 97.50%. In the case of tempotron, for Bernoulli and Poisson the calculations are very fast, taking about 7s. For Markov the time is eight times longer, while classification depends not just on the current state but also on prior states, the model must process sequences over multiple time steps, significantly increasing computational demands. Since the Poisson process does not have explicit dependencies like Markov sequences, computations remain efficient.
@@
TABLE III: Influence of supervised learning algorithm on SNN.
Type Subtype Datasets Bio-inspired Epochs $n$ Time [s] Accuracy [%] MSE MAE $R^{2}$ score
Supervised BP Bernoulli No 10 128 866.44 99.00 0.0100 0.0100 0.9600
Markov No 10 256 6211.62 100.00 0.0000 0.0000 1.0000
Poisson No 10 512 12916.37 97.50 0.0250 0.0250 0.9000
STBP Bernoulli No 10 128 2261.43 100.00 0.0000 0.0000 1.0000
Markov No 10 512 48258.21 100.00 0.0000 0.0000 1.0000
Poisson No 10 512 30093.49 96.50 0.0350 0.0350 0.8599
tempotron Bernoulli Yes 10 128 7.0 96.00 0.0400 0.0400 0.8399
Markov Yes 10 128 40.5 90.59 0.0941 0.0941 0.6234
Poisson Yes 10 128 7.23 89.50 0.1050 0.1050 0.5798
Spikeprop Bernoulli Yes 10 128 64.28 100.00 0.0000 0.0000 1.000
Markov Yes 10 64 47.0 92.34 0.0766 0.0766 0.6935
Poisson Yes 10 32 20.15 91.50 0.0850 0.0850 0.6599
Chronotron Bernoulli Yes 10 32 18.61 98.50 0.01500 0.01500 0.9400
Markov Yes 10 32 46.30 92.34 0.9300 0.0175 0.9300
Poisson Yes 10 32 113.98 90.50 0.0950 0.0950 0.6198
ReSuMe Bernoulli Yes 10 128 68.55 92.50 0.0750 0.0750 0.6999
Markov Yes 10 64 46.05 91.60 0.0840 0.0840 0.6642
Poisson Yes 10 64 21.45 91.50 0.0850 0.0850 0.6590
Table IV presents the influence of hybrid learning algorithms such as ANN-SNN conversion, STPD based on rewards, and BAL on SNN performance. For ANN-SNN conversion LZC classifier achieves high classification accuracy for all datasets. However, for Poisson process has slightly lower accuracy due to its high randomness. It is worth to stress that the Markov process requires the longest processing time due to its sequential dependencies. Moreover, the Poisson process is processed fast, despite being more stochastic. Additionally, the higher errors for Poisson data are secondary to the fact that randomness complicates the model is introduced at the boundary limit.
@@
We achieve comparable results with the reward-based and BAL learning algorithms. However in the case of reward based, for Markov and Poisson processes the network requires more neurons in layers than ANN-SNN conversion, but the computation time is shorter. For Bernoulli the network requires twice as few neurons in layers, but the computation time is five times shorter. BAL has six times longer computation time for a Poisson process than the reward based algorithm. Unlike standard Hebbian learning or STDP, BAL seems to deal with stochastic volatility more efficiently.

### v2

@@ 2.3.4 SpikeProp Algorithm
2.3.5 Chronotron Learning Rule
The Chronotron minimizes the actual and target spike times [14]
@@ 2.3.5 Chronotron Learning Rule
2.3.6 Remote Supervised Method
@@
2.4 Hybrid Learning Algorithms
Hybrid algorithms integrate elements of both supervised and unsupervised learning. These schemes combine error-driven optimization with local synaptic plasticity or bridge the gap between rate-based ANNs and event-based SNNs. Such methods aim to leverage the robustness and training efficiency of conventional deep learning while retaining the energy efficiency and temporal precision of spike-based computation. In this study we consider three representative hybrid strategies: ANN-to-SNN conversion, reward-modulated STDP, and bio-inspired active learning. Hybrid and reward-modulated learning strategies introduce controlled stochasticity through exploration mechanisms. This property allows them to maintain higher temporal variability under noisy conditions, preventing premature collapse of spike complexity and yielding more stable LZC-based representations for stochastic inputs.
2.4.1 Artificial-to-Spiking Neural Network conversion
@@
2.4.2 Reward-based STDP
Reward-based STDP modifies STDP with reinforcement learning
@@ 2.4.2 Reward-based STDP
2.4.3 Bio-inspired Active Learning
Active learning updates weights based on uncertainty [32]
@@ 3 Classification Task
The SNN neural network under consideration consists of three layers: input, hidden, and output, each containing $n$ neurons; see Figure 1. It was based on the Leaky Integrate-and-Fire (LIF) biophysical neuron model that is a fundamental component of SNNs to simulate the subthreshold behavior of the membrane potential of a neuron $U(t)\in\mathbb{R}$ [33]. Given an input vector $\mathbf{x}(t)\in\mathbb{R}^{n}$, synaptic weights $\mathbf{w}\in\mathbb{R}^{n}$, and bias $b\in\mathbb{R}$, the net input current is defined as:
@@
We considered different sets of neuron parameters and the presented results include the most efficient set of parameters for each learning algorithm. The threshold took values 1.00, 0.50, 0.40, 0.30, 0.20, 0.10, and 0.05. The decline was 0.100, 0.05, 0.03, 0.01, and the learning rates were 0.0500, 0.0100, 0.0010, 0.0098, and 0.0001. We considered neural networks comprising 16, 32, 64, 128, 256, 512, and 1024 LIF neurons per layer. The best of the results obtained are presented in Table 4, Table 5, and Table 8. A study on the effect of neuron count on transmission efficiency in the case of the Levy-Baxter neuron model has been reported in [34].
@@
where $\alpha=2$ for binary sequences. For purely random sequences, $c_{\alpha}(\textbf{x}_{n}^{1})$ converges to 1, while for deterministic sequences, it approaches 0. This measure effectively approximates the entropy of ergodic stochastic processes, making it a valuable tool to evaluate randomness and structure within the output of the neural network [25, 35]. For completeness, we performed our analysis for information sources for which the subsequent symbols are more and more correlated [36].
4 Input Datasets
In this study, we employed synthetic binary datasets designed to represent different types of temporal dependencies and levels of stochasticity. We consider three types of dataset that consist of Bernoulli sequences (uncorrelated), two-state Markov, and Poisson processes [37]. Poisson spike trains have been widely used to model neural activity since early studies, owing to their biologically realistic temporal dynamics and the characteristic proportionality between the mean and variance of neural responses, which gives rise to variable interspike intervals [38]. Experimental data have confirmed this Poisson property [38, 37]. The binary sequence of each process is converted into inputs to LIF models with specified $\Delta t=1ms$ [38]. For the evaluation of the proposed classification task, we applied accuracy the mean squared error (MSE) and $R^{2}$ score.
Bernoulli sequences are generated from independent Bernoulli processes, where each binary element $b_{i}$ in a given sequence is drawn independently with probability p. For our experiments, we generate two sets of binary sequences $B_{1}$ and $B_{2}$, each sequence of length 1024, using different values of $p$. The independence of events allows us to assess the system’s performance under random binary outcomes with varying probabilities.
Markov sequences are generated from a first-order Markov process, where the probability of each element $s_{i}$ in a given sequence depends on the previous state $s_{i-1}$. For our experiments, two sets of binary sequences of length 1024, $M_{1}=\{m_{1,k}\},k=1,2,\dots 1000$ and $M_{2}=\{m_{2,k}\},k=1,2,\dots 1000$ are generated with transition probabilities defining the dependency structure. We explore different transition probability configurations to evaluate the impact of state dependencies on classification accuracy.
Poisson sequences are generated from a Poisson process, where events occur independently at a constant average rate $\lambda$. For each Poisson process, spike trains are generated with rate parameters $\lambda_{x}$ and $\lambda_{y}$. These sequences are tested under various rate configurations to observe the effect of spike timing variability on system behavior and classification performance. This variety of data we employed allows us to systematically assess the strengths and limitations of classical and bio-inspired learning algorithms under different input regimes. Bernoulli sequences test the ability to classify memoryless inputs, Markov sequences require short-term memory handling, and Poisson sequences emulate real-world neural signal variability. Using synthetic data also enables precise control over statistical properties and eliminates noise or artifacts that can confound real-world datasets. This makes it possible to isolate and compare the intrinsic capabilities of the learning rules. Performance was evaluated using classification accuracy, mean squared error (MSE), and $R^{2}$ score.
This synthetic design allows precise control over the temporal structure and stochasticity of inputs, which is critical when evaluating biologically inspired learning rules under known statistical regimes. Unlike benchmark datasets, our generative setup isolates key dynamics relevant to spiking computation.
In addition to the synthetic temporal datasets, we also evaluated the proposed approach on binarized versions of two widely used benchmark datasets: MNIST and N-MNIST, restricted to the digits 0 and 1. The MNIST01 subset consists of grayscale handwritten digits from the original MNIST collection, filtered to include only classes 0 and 1, resulting in a clean two-class pattern-recognition task. Each image was normalized and flattened into a 1,024-dimensional input vector (32 $\times$ 32 cropping/reshaping), consistent with the dimensionality used in our spike-based models. This reduced task serves as a controlled setting to isolate the behavior of the SNN learning rules without confounding factors such as multi-class overlap or high intra-class variability. The N-MNIST01 subset follows the same class restriction but is derived from the neuromorphic N-MNIST dataset, where event-based recordings of moving MNIST digits are captured using a Dynamic Vision Sensor (DVS). Events corresponding to digits 0 and 1 were retained, and the resulting spike streams were temporally discretized into the same fixed-length input representation used for all other experiments. This allows direct comparison between conventional image-based inputs (MNIST01) and event-driven neuromorphic inputs (N-MNIST01) under identical network architectures and training procedures. Input characteristics for the MNIST01 and N-MNIST01 subsets used in our experiments are shown in Table 3. Both MNIST01 and N-MNIST01 contain only the digits 0 and 1, enabling a controlled binary classification setting. For M-NIST01, images are normalized and converted to a fixed 1,024-dimensional vector. For N-MNIST01, DVS events corresponding to digits 0 and 1 are aggregated into 32 $\times$ 32 bins, producing spike-based representations with identical dimensionality, allowing direct comparison across static and event-driven modalities.
Table 3: Input characteristics for the MNIST01 and N-MNIST01 subsets used in our experiments.
Dataset Samples (train/test) Input type Dim.
MNIST01 12,665 / 2,115 Static grayscale images 1024
N-MNIST01 12,665 / 2,115 Event-based spike streams 1024
The restriction to a binary (0/1) setting is intentional and serves to isolate the effect of learning rules on temporal representations under controlled conditions, rather than to compete with multi-class benchmark performance.
5 Results and Discussion
In contrast to conventional evaluations that focus primarily on classification accuracy, our analysis emphasizes how different learning paradigms reshape the temporal information structure of spike trains. Lempel-Ziv complexity is therefore treated as a primary descriptor, while accuracy, MSE, and $R^{2}$ scores are interpreted as secondary outcomes resulting from these underlying structural changes.
We consider classical and bio-inspired learning algorithms in SNNs. Our prior work [1] introduced LZC as an analysis tool for assessing the impact of neuron models on SNN behavior. Here, we move beyond descriptive use and evaluate how different learning paradigms reshape class-conditional complexity profiles when LZC is used as a decision-relevant temporal descriptor within the classification pipeline. The focus is on characterizing learning-rule-induced changes in temporal organization and relating them to accuracy-efficiency trade-offs under controlled input statistics.
@@ 5 Results and Discussion
Table 4 presents the influence of the unsupervised learning algorithm such as Hebbian, STDP, and SDSP on the performance of SNNs. All computations were performed for 10 epochs on an Intel(R) Core(TM) i7-14700F CPU @ 2.10 GHz. A general finding of our study is the pronounced dependence of both classification accuracy and computational overhead on the choice of learning algorithm and the underlying characteristics of the input data. Observed classification accuracies ranged from 89.50% to 100.00% across the evaluated methods, whereas computational times exhibited substantial variation, ranging from approximately 7 seconds to over 48,000 seconds.
@@
Table 4 presents the influence of the unsupervised learning algorithm such as Hebbian, STDP, and SDSP on the performance of SNNs. The results obtained for the Hebbian algorithm demonstrate that, within the classical framework, high classification accuracy was achieved for the Bernoulli and Markov processes, with values of 99.00% and 95.79%, respectively. These results were accompanied by relatively long computation times of 52.93 s and 74.66 s. Furthermore, MSE values of 0.0011 and 0.00422, suggest minimal prediction errors. The coefficient of determination $R^{2}$ scores of 0.9600 and 0.8313 further confirm a strong agreement between predicted and actual values. For the Poisson process, a classification accuracy of 90.00% was observed, with a $R^{2}$ score of 0.9074. The corresponding MSE values were 0.1000, while the $R^{2}$ score of 0.5998 indicates a moderate correlation between the predictions and the actual values. It is worth emphasizing that the accuracy for the Poisson process is achieved for a twice smaller number of neurons in the layers (for Bernoulli and Markov processes, we achieve similar accuracy with the same number of neurons in the layers). An increase in the number of neurons causes a decrease in accuracy below 50.00%, contrary to the case of Bernoulli and Markov processes, where accuracy increases with increasing number of neurons in layers. Thus, the worse classification results for the Poisson process may be related to the fact that events occur randomly in time, making it difficult to generalize Hebbian learning effectively. The relatively low computation time indicates that the Hebbian learning algorithm efficiently processes structured data. However, taking into account the nature of Hebbian learning it is effective when there are strong, predictable relationships between inputs and outputs, like in the case in Bernoulli or Markov processes.
@@
The introduction of other bioinspired learning algorithms, namely STDP and SDSP provides a similar level of accuracy, in a similar computational time, with the exception of the SDSP algorithm for Markov processes, where the computational time is an order of magnitude longer. However, the lower MSE for Bernoulli and Markov processes indicate that STDP fine-tunes synaptic weights more effectively than Hebbian learning, reducing classification errors.
@@
Table 4: Influence of unsupervised learning algorithms on SNN performance.
Type Subtype Dataset Bio- Ep. $n$ Time [s] Acc. [%] MSE $R^{2}$
Unsupervised Hebbian Bernoulli Yes 10 128 $51.93$ $99.00$ $0.0100$ $0.9600$
Markov Yes 10 128 $74.66$ $95.78$ $0.0422$ $0.8313$
Poisson Yes 10 64 $17.17$ $90.00$ $0.1000$ $0.5998$
STDP Bernoulli Yes 10 64 $62.08$ $99.50$ $0.0050$ $0.9800$
Markov Yes 10 128 $513.73$ $98.00$ $0.0200$ $0.9200$
Poisson Yes 10 64 $63.36$ $89.50$ $0.1050$ $0.5798$
SDSP Bernoulli Yes 10 128 $43.20$ $98.00$ $0.0200$ $0.9200$
Markov Yes 10 128 $46.10$ $93.87$ $0.0613$ $0.7549$
Poisson Yes 10 64 $17.83$ $92.50$ $0.0750$ $0.6999$
Table 5 presents the influence of the supervised learning algorithms such as backpropagation, tempotron, SpikeProp, Chronotron and ReSuMe on the performance of SNNs. The BP algorithm achieves perfect classification accuracy for Bernoulli, Markov, and Poisson processes, as evidenced by zero MSE, indicating no classification errors. The $R^{2}$ score of 1.000 further confirms its ability to perfectly predict sequence outputs. However, this unmatched accuracy comes at a significant computational cost, with processing times, making BP the most resource-intensive supervised method.
Although BP is theoretically optimal in terms of predictive performance, its high computational demands render it impractical for real-time or large-scale applications where efficiency is crucial. This trend also holds for the STBP algorithm, but with a significantly greater time expenditure than BP. In turn, in [29], the authors state that they studied Bernoulli sources; however, in their network, the input layer consisted of LIF neurons, while the remaining layers comprised perceptrons trained using the tempotron algorithm. They did not present results for this datasets. We achieve high classification accuracy, reaching 99.00% for Bernoulli process and 100.00% for Markov process, while in the case of Poisson process the level of accuracy in lower, namely 97.50%. In the case of tempotron, for Bernoulli and Poisson the calculations are very fast, taking about 7s. For Markov the time is eight times longer, while classification depends not just on the current state but also on prior states, the model must process sequences over multiple time steps, significantly increasing computational demands. Since the Poisson process does not have explicit dependencies like Markov sequences, computations remain efficient.
@@
Table 5: Influence of supervised learning algorithms on SNN performance.
Type Subtype Dataset Bio- Ep. $n$ Time [s] Acc. [%] MSE $R^{2}$
Supervised BP Bernoulli No 10 128 $866.44$ $99.00$ $0.0100$ $0.9600$
Markov No 10 256 $6211.62$ $100.00$ $0.0000$ $1.0000$
Poisson No 10 512 12 916.3712\,916.37 $97.50$ $0.0250$ $0.9000$
MNIST No 10 128 $207.91$ $99.05$ $0.0094$ $0.9619$
Fashion MNIST No 10 128 $1106.38$ $95.75$ $0.0425$ $0.8300$
Fashion MNIST No 10 128 $203.76$ $95.64$ $0.0435$ $0.8257$
N-MNIST No 10 128 $1508.40$ $99.23$ $0.0072$ $0.9715$
DVS Gesture No 10 128 $255.77$ $100.00$ $0.0000$ $1.0000$
SHD No 10 128 $384.71$ $95.26$ $0.0474$ $0.8103$
STBP Bernoulli No 10 128 $2261.43$ $100.00$ $0.0000$ $1.0000$
Markov No 10 512 48 258.2148\,258.21 $100.00$ $0.0000$ $1.0000$
Poisson No 10 512 30 093.4930\,093.49 $96.50$ $0.0350$ $0.8599$
MNIST No 10 128 $885.10$ $99.83$ $0.0140$ $0.9946$
Fashion MNIST No 10 128 $1106.38$ $95.75$ $0.0425$ $0.9558$
N-MNIST No 10 128 $1061.43$ $99.63$ $0.0041$ $0.9837$
DVS Gesture No 10 128 $1401.89$ $97.92$ $0.0208$ $0.9167$
SHD No 10 128 $2229.57$ $89.57$ $0.1043$ $0.5270$
Tempotron Bernoulli Yes 10 128 $7.00$ $96.00$ $0.0400$ $0.8399$
Markov Yes 10 128 $40.50$ $90.59$ $0.0941$ $0.6234$
Poisson Yes 10 128 $7.23$ $89.50$ $0.1050$ $0.5798$
MNIST No 10 128 $32.20$ $87.55$ $0.1245$ $0.8914$
Fashion MNIST No 10 128 $38.54$ $100.00$ $0.0000$ $1.0000$
N-MNIST No 10 512 $593.90$ $98.41$ $0.1776$ $0.2865$
DVS Gesture No 10 128 $119.15$ $72.95$ $0.2710$ $0.7350$
SHD No 10 128 $68.71$ $72.04$ $0.2799$ $0.1191$
SpikeProp Bernoulli Yes 10 128 $64.28$ $100.00$ $0.0000$ $1.0000$
Markov Yes 10 64 $47.00$ $92.34$ $0.0766$ $0.6935$
Poisson Yes 10 32 $20.15$ $91.50$ $0.0850$ $0.6599$
MNIST No 10 128 $68.80$ $99.12$ $0.0129$ $0.9484$
Fashion MNIST No 10 128 $92.49$ $93.68$ $0.0632$ $0.7471$
N-MNIST No 10 128 $383.30$ $98.38$ $0.0162$ $0.9348$
DVS Gesture No 10 128 $166.38$ $100.00$ $0.0000$ $1.0000$
SHD No 10 128 $99.37$ $91.94$ $0.0806$ $0.6775$
Chronotron Bernoulli Yes 10 32 $18.61$ $98.50$ $0.0150$ $0.9400$
Markov Yes 10 32 $46.30$ $92.34$ $0.9300$ $0.0175$
Poisson Yes 10 32 $113.98$ $90.50$ $0.0950$ $0.6198$
MNIST No 10 256 16 329.1016\,329.10 $76.18$ $0.2382$ $0.0432$
Fashion MNIST No 10 128 $2394.12$ $75.25$ $0.2475$ $0.0100$
N-MNIST No 10 256 10 299.0710\,299.07 $75.41$ $0.1373$ $0.4482$
DVS Gesture No 10 128 $197.17$ $87.50$ $0.1250$ $0.8980$
SHD No 10 64 $100.18$ $81.99$ $0.1801$ $0.2792$
ReSuMe Bernoulli Yes 10 128 $68.55$ $92.50$ $0.0750$ $0.6999$
Markov Yes 10 64 $46.05$ $91.60$ $0.0840$ $0.6642$
Poisson Yes 10 64 $21.45$ $91.50$ $0.0850$ $0.6590$
MNIST No 10 128 $56.04$ $98.55$ $0.0145$ $0.9416$
Fashion MNIST No 10 128 $63.34$ $95.07$ $0.0493$ $0.8029$
N-MNIST No 10 128 $438.12$ $98.68$ $0.0132$ $0.9470$
DVS Gesture No 10 128 $171.14$ $100.00$ $0.0000$ $1.0000$
SHD No 10 128 $101.25$ $91.94$ $0.0806$ $0.6775$
@@
Table 8 presents the influence of hybrid learning algorithms such as ANN-SNN conversion, STPD based on rewards, and BAL on SNN performance. For ANN-SNN conversion, the LZC-based classifier achieves high classification accuracy across all datasets. However, performance for Poisson-distributed inputs remains lower than for Bernoulli and Markov sources due to higher temporal variability. We achieve comparable results with the reward-based and BAL learning algorithms. However in the case of reward based, for Markov and Poisson processes the network requires more neurons in layers than ANN-SNN conversion, but the computation time is shorter. For Bernoulli the network requires twice as few neurons in layers, but the computation time is five times shorter. BAL has six times longer computation time for a Poisson process than the reward based algorithm. Unlike standard Hebbian learning or STDP, BAL seems to deal with stochastic volatility more efficiently.